# Thermodynamic topology of Black Holes in $F(R)$-Euler-Heisenberg gravity’s Rainbow

### v1

@@ VII Conclusion
The study of black hole thermodynamics delves into the intriguing connections between thermodynamic properties and the topological features of black holes. This field often involves examining critical points in black hole phase diagrams and assigning topological charges to these points. A notable approach is Duan’s topological current $\phi$-mapping theory, which introduces the concept of topological charges to these critical points. This method offers a new perspective on the thermodynamic behavior of black holes and can be applied to other types of black holes, revealing more complex topological information.
In conclusion, Our research has yielded several significant findings. We have successfully derived the field equations for $F(R)$-Euler-Heisenberg theory, establishing a framework to explore the interaction between modified gravity and non-linear electromagnetic effects. Additionally, we have obtained an analytical solution for a static, spherically symmetric, energy-dependent black hole with constant scalar curvature. Also, our analysis of black holes in F(R)-Euler-Heisenberg gravity’s Rainbow provides significant insights into their topological properties. By examining the normalized field lines and various free parameters, we identified distinct winding and total topological numbers. Our findings suggest that the parameters $(R_{0})$ and $(f_{\epsilon}=g_{\epsilon})$ influence the topological charges. The stability of these black holes was further assessed through winding numbers, highlighting the intricate relationship between these parameters and the black holes’ topological characteristics. These results are comprehensively summarized in Table I. Additionally, an overview of Tables II, III, and IV related to the photon sphere of the mentioned black hole shows that with an increase in $f_{\varepsilon}$, the permissible range of negative $\lambda$ in the first case gradually transitions into a non-permissible region in the third case. Furthermore, the QED parameter, which measures the strength of nonlinear effects, can be either positive or negative. A positive QED parameter reduces the electric field near the horizon and increases the black hole’s mass, whereas a negative QED parameter increases the electric field and decreases the mass. Based on these observations, it can be concluded that an increase in $f_{\varepsilon}$ decreases the strength of the electric field near the horizon and enhances the effects of gravity.

### v2

@@ VII Conclusion
The study of black hole thermodynamics delves into the intriguing connections between thermodynamic properties and the topological features of black holes. This field often involves examining critical points in black hole phase diagrams and assigning topological charges to these points. A notable approach is Duan’s topological current $\phi$-mapping theory, which introduces the concept of topological charges to these critical points. This method offers a new perspective on the thermodynamic behavior of black holes and can be applied to other types of black holes, revealing more complex topological information. We have successfully derived the field equations for $F(R)$-Euler-Heisenberg theory, establishing a framework to explore the interaction between modified gravity and non-linear electromagnetic effects. Additionally, we have obtained an analytical solution for a static, spherically symmetric, energy-dependent black hole with constant scalar curvature. Also, our analysis of black holes in F(R)-Euler-Heisenberg gravity’s Rainbow provides significant insights into their topological properties. By examining the normalized field lines and various free parameters, we identified distinct winding and total topological numbers. Our findings suggest that the parameters $(R_{0})$ and $(f_{\epsilon}=g_{\epsilon})$ influence the topological charges. The stability of these black holes was further assessed through winding numbers, highlighting the intricate relationship between these parameters and the black holes’ topological characteristics. These results are comprehensively summarized in Table I. In the study of the photon sphere for this model, as emphasized above, the sign of the parameter$(R_{0})$ can significantly influence the model’s behavior, transforming it into either a dS or AdS form. An interesting aspect of this model is that, in its AdS form, unlike many models where selecting different parameter values could behaviorally divide the space into black hole regions and naked singularities, this model does not exhibit a naked singularity region. It appears that, for various parameter values within the permissible range, the system can maintain its black hole behavior by displaying an unstable photon sphere. Ultimately, in the dS form, the behavior of the model’s photon sphere remains general, like other dS models studied, and does not show any particular differences.